OpenStudy (anonymous):

A challenge for the new year: If possible, find a closed form or reduction formula for the integral, \[\Large\int_0^{\pi/2}\sin^{\frac{2n+1}{2}}x\,dx\]

ganeshie8 (ganeshie8):

I'm trying something like this \[\begin{align} \int_0^{\pi/2}\sin^{\frac{2n+1}{2}}x\,dx &= \int_0^{\pi/2}(1-\cos ^2 x)^{\frac{2n-1}{4}}\cdot \sin x\,dx\\~\\ &=\int_0^{1}(1-u^2)^{\frac{2n-1}{4}}\,du\\~\\ &=\int_0^{1}\sum\limits_{k=0}^{\infty} \binom{(2n-1)/4}{k} (-u^2)^k\,du\\~\\ &=\sum\limits_{k=0}^{\infty} \binom{(2n-1)/4}{k} \frac{(-1)^k}{2k+1}\\~\\ \end{align}\] whatever that equals hmm

OpenStudy (anonymous):

Why did you make it an infinite sum?

ganeshie8 (ganeshie8):

thats because of the generalized binomial thm http://en.wikipedia.org/wiki/Binomial_series

ganeshie8 (ganeshie8):

I'm still trying to see if the sum evaluates to something that looks more like "closed form"

OpenStudy (anonymous):

i was confused all this time thinking 2n+1/2 is inside sin :/

ganeshie8 (ganeshie8):

i got a nice recurrence relation because of the symmetry after substituting away sin.. il post, one sec..

OpenStudy (anonymous):

hmm 2n+1/2 is not integer , hmm i need to think of it

ganeshie8 (ganeshie8):

For recurrence relation : \[\begin{align} I=\int_0^{\pi/2}\sin^{\frac{2n+1}{2}}x\,dx &= \int_0^{\pi/2}(1-\cos ^2 x)^{\frac{2n-1}{4}}\cdot \sin x\,dx\\~\\ &=\int_0^{1}(1-u^2)^{\frac{2n-1}{4}}\,du\\~\\ &=\frac{1}{2}\int_0^{1}t^{\frac{-1}{2}}(1-t)^{\frac{2n-1}{4}}\,dt \\~\\ &=\frac{1}{2}\int_0^{1}t^{\frac{-1}{2}}(1-t)^{q}\,dt~~~;~~ \color{Red}{\text{ (2n-1)/4}=q}\\~\\ \end{align}\] let \(I_q =\frac{1}{2}\int_0^{1}t^{\frac{-1}{2}}(1-t)^{q}\,dt \) By parts we get \[\begin{align} I_q &=(1-t)^q \frac{t^{1/2}}{1/2}\Bigg|_0^1+\frac{1}{2}\int_0^{1}\frac{t^{\frac{1}{2}}}{1/2}\cdot q(1-t)^{q-1}\,dt \\~\\ &=0+q\int_0^{1}t^{\frac{1}{2}}\cdot (1-t)^{q-1}\,dt \\~\\ &=q\left[\int_0^{1}t^{\frac{-1}{2}}\cdot (1-t)^{q-1}\,dt -\int_0^{1}t^{\frac{-1}{2}}\cdot (1-t)^{q}\,dt \right]\\~\\ &=q\left[I_{q-1} -2I_q \right]\\~\\ \end{align}\]

OpenStudy (anonymous):

haha i reached reduction formula only xD

ganeshie8 (ganeshie8):

lets see if we mess with the recurrence relation further remember gamma function ?

OpenStudy (anonymous):

yes , hmm how its involved here ?

ganeshie8 (ganeshie8):

wolfram says the integral can be represented using gamma function : \[\sum\limits_{k=0}^{\infty} \binom{(2n-1)/4}{k} \frac{(-1)^k}{2k+1} ~~=~~ \dfrac{\sqrt{\pi} ~\Gamma ((2n+3)/4)}{2~\Gamma ((2n+5)/4)}\] http://www.wolframalpha.com/input/?i=%5Csum%5Climits_%7Bk%3D0%7D%5E%7B%5Cinfty%7D+%5Cbinom%7B%282n-1%29%2F4%7D%7Bk%7D+%5Cfrac%7B%28-1%29%5Ek%7D%7B2k%2B1%7D

OpenStudy (anonymous):

ok i understand how to convert this series to gama , but where did the series come from ?

ganeshie8 (ganeshie8):

look at my first reply

OpenStudy (anonymous):

oh yes , i thught u start over >.<

OpenStudy (anonymous):

yeah thats reduce it in nice formula , now to generate gama , here is in integral of it